### v1

@@ Appendix C Proof of Theorem 1
$=(\sigma^{2}+rc)\mathsf{I}_{d}-r^{2}\mathbf{1}^{T}_{n-c-1}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c-1)\times(n-c-1)}^{-1}\!\!\!\!\!\!\!\!\!\!\!\!\mathbf{1}_{n-c-1}\mathsf{I}_{d}$ (66)
$=\left((\sigma^{2}+rc)-r^{2}\mathbf{1}^{T}_{n-c-1}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots&\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{(n-c-1)\times(n-c-1)}^{-1}\!\!\!\!\!\!\!\!\!\!\!\!\mathbf{1}_{n-c-1}\right)\mathsf{I}_{d}$ (67)
@@
Proof:  [Proof of Claim 2] For given values of $\mathbf{Y}_{j}=\mathbf{y}_{j}$, $\forall j\in[2:n-c]$, consider the variation of $\mathbf{x}_{1}$ by fixing all $\mathbf{x}_{j}$, $\forall j\in[2:n-c]$, and define,
$f(\mathbf{x}_{1})$ $=\tilde{\mathbf{\mu}}$ (70)
$=\mathbf{x}_{1}+\left(r\mathbf{1}_{n-c-1}^{T}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{n-c-1}^{-1}\otimes\mathsf{I}_{d}\right)\begin{pmatrix}\mathbf{y}_{2}-\mathbf{x}_{2}\\ \vdots\\ \mathbf{y}_{n-c}-\mathbf{x}_{n-c}\end{pmatrix}$ (71)
Then, define a new random variable $\mathbf{W}=f(\mathbf{x}_{1})+\mathbf{N}$, where $\mathbf{N}\sim\mathcal{N}\left(\mathbf{0}_{d},\tilde{\sigma}^{2}\mathsf{I}_{d}\right)$. Note that for any given values of $\mathbf{y}_{j}$, $\mathbf{x}_{j}$, $\forall j\in[2:n-c]$, $\mathbf{W}\sim\mathbf{Y}_{1}|\mathbf{Y}_{j}=\mathbf{y}_{j},j\in[2:n-c]$, (statistically equivalent). Now, consider the $(\epsilon,\delta)$-DP constraint in (54).
$\mathbb{P}(M(\mathbf{x}_{1})\in\mathcal{A}|$ $M(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[2:n-c])\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{1}^{\prime})\in\mathcal{A}|M(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[2:n-c])+\delta$ (72)
@@
$f^{\prime}(\hat{\sigma})$ $=\frac{1}{2\pi}\left(-\frac{1}{\sigma^{2}}-\frac{\epsilon}{2}\right)e^{-\frac{1}{2}\left(\frac{1}{\hat{\sigma}^{2}}+\frac{\epsilon^{2}\hat{\sigma}^{2}}{4}-\epsilon\right)}-\frac{e^{\epsilon}}{2\pi}\left(\frac{1}{\sigma^{2}}-\frac{\epsilon}{2}\right)e^{-\frac{1}{2}\left(\frac{1}{\hat{\sigma}^{2}}+\frac{\epsilon^{2}\hat{\sigma}^{2}}{4}+\epsilon\right)}$ (78)
@@
$\left(\sigma^{2}+rc-r^{2}\mathbf{1}_{n-c-1}^{T}\begin{pmatrix}\sigma^{2}+rc&r&\dotsc&r\\ r&\sigma^{2}+rc&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}+rc\end{pmatrix}_{n-c-1}^{-1}\mathbf{1}_{n-c-1}\right)$ $\geq\sigma^{2}_{\epsilon,\delta}$ (80)
$\sigma^{2}+rc-r^{2}\mathbf{1}_{n-c-1}^{T}\big{(}(\sigma^{2}+r(c-1))\mathsf{I}_{n-c-1}+r\mathbf{1}_{n-c-1}\mathbf{1}_{n-c-1}^{T}\big{)}^{-1}\mathbf{1}_{n-c-1}$ $\geq\sigma^{2}_{\epsilon,\delta}$ (81)
@@
$\sigma^{2}+rc-r^{2}\mathbf{1}^{T}_{n-c-1}\left(\frac{1}{\sigma^{2}+r(c-1)}I_{n-c-1}-\frac{\frac{r}{(\sigma^{2}+r(c-1))^{2}}\mathbf{1}_{n-c-1}\mathbf{1}_{n-c-1}^{T}}{1+\frac{r}{\sigma^{2}+r(c-1)}\mathbf{1}_{n-c-1}^{T}\mathbf{1}_{n-c-1}}\right)\mathbf{1}_{n-1}$ $\geq\sigma^{2}_{\epsilon,\delta}$ (82)
@@
$\geq\sigma^{2}+\left(\frac{(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2(n-1)}-\frac{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)}\right)(n-1)$ (86)
$=\frac{1}{2}\left(2\sigma^{2}+(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}\right)$ (87)
@@
$\sigma^{2}-r$ $\geq\sigma^{2}-\frac{(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2(n-1)}-\frac{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)}$ (91)
@@
$\sigma^{2}+r(n-1)\geq\sigma^{2}-\frac{\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n-1)}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)}=\frac{\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)}>0.$ (93)
@@
$\geq\sigma^{2}+\frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(c-1)}+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(c-1)}$ (94)
$\geq\frac{-((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)-2\sigma^{2}+\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(c-1)}$ (95)
@@
$p^{2}-|q|^{2}$ $=((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})$
$\quad-((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}-4\sigma^{2}((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)-4\sigma^{4}$ (98)
@@
For any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$, the optimum $r$ that satisfies $(\epsilon,\delta)$-DP is given by,
$r_{*}$ $=\arg\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ (101)
=$=\begin{cases}a-b,&c=0\\ \frac{-\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(n-1)(c-1)}+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)(c-1)},&c>1\end{cases}$ (102)
@@
For the case where $t=n$, $\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ is a decreasing function in $\sigma^{2}$ and,
$\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ $=\lim_{\sigma^{2}\to\infty}\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},r,\bm{\alpha})$ (105)
=$=\begin{cases}O\left(\frac{d\sigma^{2}_{\epsilon,\delta}}{(n-c)^{2}}\right),&\text{if }(n-c)^{2}\gg d\\ O(1),&otherwise.\end{cases}$ (106)
@@
Proof:  For any given $\sigma^{2}$, $r$, $\mathcal{F}$ and $\mathcal{C}$, from Proposition 1,
$\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r(|\mathcal{F}|-1)}\right)^{-1}$ (107)
@@
For the case of no dropouts, i.e., $t=n$, $\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})$ is a decreasing function in $\sigma^{2}$ for any $\mathcal{C}$ satisfying $0\leq|\mathcal{C}|<n$.
@@
$\min_{r}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},r,\bm{\alpha})=\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r_{*}(n-1)}\right)^{-1}$ (108)
as $|\mathcal{F}|=n$. Let $f(\sigma^{2})=\sigma^{2}+r_{*}(n-1)$. We prove that (108) decreases with $\sigma^{2}$ by showing that $f(\sigma^{2})$ is a decreasing function in $\sigma^{2}$ for each of the three cases in (102).
@@
$=1+\frac{n-2}{2}-\frac{(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n-2)^{2}+2(n-1)(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}$ (110)
@@
$\quad-(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}(n-2)^{4}-4(n-1)^{2}(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}-4(n-1)(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})$ (112)
$=4(n-1)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}(n-2)^{2}+4(n-1)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n^{2}\sigma^{2}-(n-2)^{2}(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta}))$
@@
$=1+(n-1)\left(\frac{(\sigma^{2}_{\epsilon,\delta}-2\sigma^{2})(\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2))+(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{(\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2))^{2}}\right)$ (119)
@@
$r*$ $=\frac{-(n\!-\!2)(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})\!-\!\sigma^{2}c}{2(n\!-\!1)(c\!-\!1)}\!+\!\frac{\sqrt{((n\!-\!2)(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})\!-\!\sigma^{2}c)^{2}+4(n\!-\!c\!-\!1)\sigma^{2}(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})}}{2(n\!-\!1)(c\!-\!1)}$ (121)
@@
$=\frac{\sigma^{2}(c\!-\!n)\!+\!\sigma^{2}_{\epsilon,\delta}(n\!-\!2)}{2(c\!-\!1)}\!+\!\frac{\sqrt{((n\!-\!2)(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})-c\sigma^{2})^{2}\!+\!4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(c-1)}$ (123)
$\frac{df(\sigma^{2})}{d\sigma^{2}}$ $=\frac{c-n}{2(c-1)}+\frac{\sigma^{2}(n-c)^{2}-\sigma^{2}_{\epsilon,\delta}(n^{2}-nc-2n+2)}{2(c-1)\sqrt{((n\!-\!2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})\!-\!c\sigma^{2})^{2}+4(n\!-\!c\!\!-\!1)\sigma^{2}(\sigma^{2}\!-\!\!\sigma^{2}_{\epsilon,\delta})}}$ (124)

### v2

@@ Appendix C Proof of Theorem 1
$=(\sigma^{2}+rc)\mathsf{I}_{d}$
$\quad-r^{2}\mathbf{1}^{T}_{n\!-\!c\!-\!1}\!\!\begin{pmatrix}\sigma^{2}\!+\!rc\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ r\!\!\!&\!\!\!\sigma^{2}\!+\!rc\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ \vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\\ r\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!\sigma^{2}\!+\!rc\end{pmatrix}^{\!-1}\!\!\!\!\!\mathbf{1}_{n\!-\!c\!-\!1}\otimes\mathsf{I}_{d}$ (88)
$\!=\!\!\left(\!\!(\sigma^{2}\!+\!rc)\!-\!r^{2}\mathbf{1}^{T}_{n\!-\!c\!-\!1}\!\!\begin{pmatrix}\sigma^{2}+rc\!\!\!\!&\!\!\!\!r\!\!\!\!&\!\!\!\!\dotsc\!\!\!\!&\!\!\!\!r\\ r\!\!\!\!&\!\!\!\!\sigma^{2}+rc\!\!\!\!&\!\!\!\!\dotsc\!\!\!\!&\!\!\!\!r\\ \vdots\!\!\!\!&\!\!\!\!\vdots\!\!\!\!&\!\!\!\!\vdots\!\!\!\!&\!\!\!\!\vdots\!\!\!\!&\!\!\!\!\\ r\!\!\!\!&\!\!\!\!r\!\!\!\!&\!\!\!\!\dotsc\!\!\!\!&\!\!\!\!\sigma^{2}+rc\!\!\!\!\!\end{pmatrix}^{-1}\!\!\!\!\!\!\!\!\!\mathbf{1}_{n\!-\!c\!-\!1}\!\!\!\right)\!\!\mathsf{I}_{d}$ (89)
@@
Proof:  [Proof of Claim 2] For given values of $\mathbf{Y}_{j}=\mathbf{y}_{j}$, $\forall j\in[2:n-c]$, consider the variation of $\mathbf{x}_{1}$ by fixing all $\mathbf{x}_{j}$, $\forall j\in[2:n-c]$, and define $f(\mathbf{x}_{1})=\tilde{\mathbf{\mu}}$. Then, define a new random variable $\mathbf{W}=f(\mathbf{x}_{1})+\mathbf{N}$, where $\mathbf{N}\sim\mathcal{N}\left(\mathbf{0}_{d},\tilde{\sigma}^{2}\mathsf{I}_{d}\right)$. Note that for any given values of $\mathbf{y}_{j}$, $\mathbf{x}_{j}$, $\forall j\in[2:n-c]$, $\mathbf{W}\sim\mathbf{Y}_{1}|\mathbf{Y}_{j}=\mathbf{y}_{j},j\in[2:n-c]$, (statistically equivalent). Now, consider the $(\epsilon,\delta)$-DP constraint in (220).
$\mathbb{P}(\hat{M}(\mathbf{x}_{1})\in\mathcal{A}|\hat{M}(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[2:n-c])$
$\quad\leq e^{\epsilon}\mathbb{P}(\hat{M}(\mathbf{x}_{1}^{\prime})\in\mathcal{A}|\hat{M}(\mathbf{x}_{j})=\mathbf{y}_{j},j\in[2:n-c])+\delta$ (92)
@@
$f(\hat{\sigma})$ $=\Phi\left(\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\left(-\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)$ (97)
@@
$f^{\prime}(\hat{\sigma})$ $=\frac{1}{2\pi}\left(-\frac{1}{\sigma^{2}}-\frac{\epsilon}{2}\right)e^{-\frac{1}{2}\left(\frac{1}{\hat{\sigma}^{2}}+\frac{\epsilon^{2}\hat{\sigma}^{2}}{4}-\epsilon\right)}$
$\quad-\frac{e^{\epsilon}}{2\pi}\left(\frac{1}{\sigma^{2}}-\frac{\epsilon}{2}\right)e^{-\frac{1}{2}\left(\frac{1}{\hat{\sigma}^{2}}+\frac{\epsilon^{2}\hat{\sigma}^{2}}{4}+\epsilon\right)}$ (99)
@@
$\left(\!\!\sigma^{2}\!+\!rc\!-\!r^{2}\mathbf{1}_{n\!-\!c\!-\!1}^{T}\!\!\begin{pmatrix}\sigma^{2}+rc\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ r\!\!\!&\!\!\!\sigma^{2}+rc\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!r\\ \vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\!\!\!&\!\!\!\vdots\\ r\!\!\!&\!\!\!r\!\!\!&\!\!\!\dotsc\!\!\!&\!\!\!\sigma^{2}+rc\!\!\end{pmatrix}^{-1}\!\!\!\!\!\!\!\!\mathbf{1}_{n-c-1}\!\!\right)\!\geq\!\sigma^{2}_{\epsilon,\delta}$ (101)
$\sigma^{2}+rc-r^{2}\mathbf{1}_{n-c-1}^{T}\left((\sigma^{2}+r(c-1))\mathsf{I}_{n-c-1}\right.$
$\left.\qquad\qquad+r\mathbf{1}_{n-c-1}\mathbf{1}_{n-c-1}^{T}\right)^{-1}\mathbf{1}_{n-c-1}\geq\sigma^{2}_{\epsilon,\delta}$ (102)
@@
$\sigma^{2}+rc-r^{2}\mathbf{1}^{T}_{n-c-1}\left(\frac{1}{\sigma^{2}+r(c-1)}I_{n-c-1}\right.$
$\left.\quad-\frac{\frac{r}{(\sigma^{2}+r(c-1))^{2}}\mathbf{1}_{n-c-1}\mathbf{1}_{n-c-1}^{T}}{1+\frac{r}{\sigma^{2}+r(c-1)}\mathbf{1}_{n-c-1}^{T}\mathbf{1}_{n-c-1}}\right)\mathbf{1}_{n-1}\geq\sigma^{2}_{\epsilon,\delta}$ (103)
@@
$\geq\sigma^{2}+\left(\frac{(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2(n-1)}\right.$
$\left.-\frac{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)}\right)(n-1)$ (107)
$=\frac{1}{2}\left(2\sigma^{2}+(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})\right.$
$\left.\quad-\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}\right)$ (108)
@@
$\sigma^{2}-r$ $\geq\sigma^{2}-\frac{(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2(n-1)}$
$-\frac{\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)}$ (112)
@@
$\sigma^{2}+r(n-1)$ $\geq\sigma^{2}-\frac{\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n-1)}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)}$ (114)
$=\frac{\sigma^{2}\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)}>0.$ (115)
@@
$\geq\sigma^{2}+\frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(c-1)}$
$+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(c-1)}$ (116)
$\geq\frac{-((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)-2\sigma^{2}}{2(c-1)}$
$+\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(c-1)}$ (117)
@@
$p^{2}-|q|^{2}$
$=((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})$
$\quad-((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}$
$\quad-4\sigma^{2}((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)-4\sigma^{4}$ (120)
@@
For any fixed $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$, the optimum $r=\rho\sigma^{2}$ that satisfies $(\epsilon,\delta)$-DP is given by,
$r_{*}$ $=\arg\min_{r}\min_{\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}_{all}})$ (123)
=$=\begin{cases}a-b,&c=0\\ \frac{-\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c}{2(n-1)(c-1)}\\ +\frac{\sqrt{((n-2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})-\sigma^{2}c)^{2}+4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(n-1)(c-1)},&c>1\end{cases}$ (124)
@@
For the case where $t=n$, $\min_{r}\min_{\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}_{all}})$ is a decreasing function in $\sigma^{2}$ and,
$\min_{\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}}\min_{r}$ $\min_{\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}_{all}})$
$=\lim_{\sigma^{2}\to\infty}\min_{r}\min_{\boldsymbol{\alpha}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}_{all}})$ (127)
=$=\begin{cases}O\left(\frac{d\sigma^{2}_{\epsilon,\delta}}{n(n-c)}\right),&\text{if }n(n-c)\gg d\\ O(1),&otherwise.\end{cases}$ (128)
@@
Proof:  For any given $\sigma^{2}$, $r$, $\mathcal{U}$, from Proposition 1,
$\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(n,\mathcal{U},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}})$ $=\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}+r(|\mathcal{U}|-1)}\right)^{-1}$ (129)
@@
For the case of no dropouts, i.e., $t=n$, $\min_{r}\min_{\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha}_{\mathcal{U}_{all}})$ is a decreasing function in $\sigma^{2}$ for any $\mathcal{U}_{col}$ satisfying $0\leq|\mathcal{U}_{col}|<n$.
@@
$\min_{r}\min_{\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(n,\mathcal{U}_{all},\sigma^{2},r,\boldsymbol{\alpha})=\left(1+\frac{n/d}{\sigma^{2}+r_{*}(n-1)}\right)^{-1}$ (130)
as $|\mathcal{U}|=n$. Let $f(\sigma^{2})=\sigma^{2}+r_{*}(n-1)$. We prove that (130) decreases with $\sigma^{2}$ by showing that $f(\sigma^{2})$ is a decreasing function in $\sigma^{2}$ for each of the three cases in (124).
@@
$=1+\frac{n-2}{2}$
$-\frac{(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n-2)^{2}+2(n-1)(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{2\sqrt{(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}+4(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}$ (132)
@@
$\quad-(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}(n-2)^{4}-4(n-1)^{2}(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}$
$\quad-4(n-1)(n-2)^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta})$ (134)
$=4(n-1)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})^{2}(n-2)^{2}$
$\quad+4(n-1)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})(n^{2}\sigma^{2}-(n-2)^{2}(2\sigma^{2}-\sigma^{2}_{\epsilon,\delta}))$
@@
$\!=\!1\!+\!(n\!-\!1)\left(\frac{(\sigma^{2}_{\epsilon,\delta}-2\sigma^{2})(\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2))}{(\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2))^{2}}\right.$
$\left.\qquad+\frac{(n-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}{(\sigma^{2}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2))^{2}}\right)$ (141)
@@
$r*=\frac{-(n\!-\!2)(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})\!-\!\sigma^{2}c}{2(n\!-\!1)(c\!-\!1)}$
$+\frac{\sqrt{((n\!-\!2)(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})\!-\!\sigma^{2}c)^{2}+4(n\!-\!c\!-\!1)\sigma^{2}(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})}}{2(n\!-\!1)(c\!-\!1)}$ (143)
@@
$=\frac{\sigma^{2}(c\!-\!n)\!+\!\sigma^{2}_{\epsilon,\delta}(n\!-\!2)}{2(c\!-\!1)}$
$+\frac{\sqrt{((n\!-\!2)(\sigma^{2}\!-\!\sigma^{2}_{\epsilon,\delta})-c\sigma^{2})^{2}\!+\!4(n-c-1)\sigma^{2}(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})}}{2(c-1)}$ (145)
$\frac{df(\sigma^{2})}{d\sigma^{2}}=\frac{c-n}{2(c-1)}$
$+\frac{\sigma^{2}(n-c)^{2}-\sigma^{2}_{\epsilon,\delta}(n^{2}-nc-2n+2)}{2\!(c\!-\!1)\!\sqrt{((n\!-\!2)(\sigma^{2}-\sigma^{2}_{\epsilon,\delta})\!-\!c\sigma^{2})^{2}\!+\!4(n\!-\!c\!\!-\!1)\sigma^{2}(\sigma^{2}\!-\!\!\sigma^{2}_{\epsilon,\delta})}}$ (146)